the coffee machine in the pi cafe has three dispensers p,q, and r. the probabilities that each malfunctions on a given day are 0.007,0.003 and 0.001 respectively.find the probability that exact one of the dispensers malfunctions on a given day.
1
Expert's answer
2020-04-13T11:01:21-0400

From the given data probability of malfunction is given by,

"P(p)=0.007\\
P(q)=0.003\\
P(r)=0.001\\"

And for each of the machines , probability of not malfunctioning is given by,

"P'(p)=1-0.007=0.993\\
P'(q)=1-0.003=0.997\\
P'(r)=1-0.001=0.999\\"


Since malfunctions are independent,

Probability that exact one of the dispensers malfunctions is given by,


"P(E)=P'(p)P'(q)P(r)+P'(p)P(q)P'(r)+P(p)P'(q)P'(r)\\
P(E)=0.993\times0.997\times0.001+0.993\times0.003\times0.999+0.007\times0.997\times0.999\\
P(E)=0.00099+0.00297+0.00697\\
P(E)=0.0109"